Karissa is having 14th party and her parents are willing for spend $150. Each guest $8 per head and Decorations cost $63. Write an equation and solve with workings

Answer by ewatrrr(24785)

8x + 63 = 150
8x = 87
x = 87/8 = 10 7/8 guests :)
10 guests to stay under budget